OpenStudy (anonymous):

The domain of the function f(x) = 3x3 is {2, 5} . what is the range of the funtion ? The range of the function f(x) = 3x3 when its domain is {2, 5} is ...

OpenStudy (michele_laino):

Please, is your function, this? \[f(x)=3x ^{3}\]

OpenStudy (anonymous):

I edit it @Michele_Laino , its up there

OpenStudy (michele_laino):

Ok! your functionis continue, it's an increasing function and monotonic, so the range is the subsequent close interval of the real line: [3*2^3, 3*5^3]=[24,375]

OpenStudy (michele_laino):

@sunnyday88 

OpenStudy (anonymous):

Im sorry im really confused . so 24,375 is the range ?

OpenStudy (michele_laino):

It's the image of the closed interval [2,5]. Nevertheless, as function, we can write: \[x \rightarrow f(x)=3x ^{3}:\mathbb{R} \rightarrow \mathbb{R} \]

OpenStudy (anonymous):

okay thankss

OpenStudy (michele_laino):

thanks!